During a head-on car crash, a passenger in the front seat of the car accelerates from 13.0 m/s to rest in 0.10 seconds. The passenger is holding a 20.9 kg child on her lap. The passenger is wearing a seat belt, but the child is not. What force must the passenger exert on the child to prevent the child from smashing into the dashboard or windshield?


1
Expert's answer
2021-01-07T10:56:39-0500

During the incendent the passenger (and the child as well) experienced the following force:


"F = ma"

where "m = 20.9kg" is the mass of the child, and "a" is her acceleration. By definition, the acceleration is:


"a = \dfrac{v}{t}"

where "v = 13m\/s" is her initial speed, and "t = 0.1s" is the time she comes to rest. The force the the passenger must exert on the child is equal and opposite to the force "F". Thus, obtain:


"F = \dfrac{mv}{t}\\
F = \dfrac{20.9\cdot 13}{0.1} = 2717N"

Answer. 2717 N.